OpenStudy (bananas1234):

V varies inversely with T and V = 12 when T = 6. Which equation shows this relationship? I got V=72T

Nnesha (nnesha):

nope "inverse variation" not direct

Nnesha (nnesha):

which equation did u use to find k ??

OpenStudy (bananas1234):

k = xy

OpenStudy (bananas1234):

so wouldn't that be k = 12 times 6 and that = 72.

Nnesha (nnesha):

hmm okay so replace k by 72 and then solve for V(y)

Nnesha (nnesha):

yep k value is right but that not suppose to be v = 72t remember inverse variation is y = k/x

OpenStudy (bananas1234):

ok so V(y) = 72/x

Nnesha (nnesha):

yep right but remember in this question x is t

OpenStudy (bananas1234):

ohh ok

Nnesha (nnesha):

inverse variation \[\huge\rm y = \frac{ k }{ x }\] \[\huge\rm y = kx \] direct variation

OpenStudy (bananas1234):

so V = 72/t ?

Nnesha (nnesha):

yep